OpenStudy (anonymous):

Jessica is baking a cake. The recipe says that she has to mix 96 grams of sugar to the flour. Jessica knows that 1 cup of this particular sugar has a mass of 128 grams. She added one over two of a cup of sugar to the flour. Should Jessica add more sugar to make the exact recipe or did she go over and by what amount? She went over by two over three of a cup She needs to add two over three of a cup She went over by one over four of a cup She needs to add one over four of a cup

OpenStudy (anonymous):

what is the modual lession hun

OpenStudy (anonymous):

01.08 Module One Review and Practice Test :3

OpenStudy (anonymous):

d bbe

OpenStudy (anonymous):

thx <3 ohmergod these questions are long o.o

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

Nancy wants to make a graph to show the relationship between the temperature of water, in degrees Celsius, and the time, in seconds, for which water is heated. She plots the following points: (0, 50), (1, 60), (2, 70), (3, 80), (4, 90) She uses the following steps to plot the graph: Label origin as (0, 0) On x-axis label Time in minutes. On y-axis label Temperature in degrees Celsius. The scale on x-axis starts from 0 and goes up to 5 at intervals of 1. The scale on the y-axis starts from 0 and goes up to 200 at intervals of 50. Which of the following best describes the graph? It will not be spread out horizontally across the entire coordinate plane because in Step 4 Nancy selected an incorrect scale on the axis. It will not be spread out vertically across the entire coordinate plane because in Step 5 Nancy selected an incorrect scale on the y-axis. It will not be spread out horizontally across the entire coordinate plane because in Step 2 Nancy plotted time instead of temperature on the axis. It will not be spread out vertically across the entire coordinate plane because in Step 3 Nancy plotted temperature instead of time on the y-axis.

OpenStudy (anonymous):

the question about x*x=blank it equals x2

OpenStudy (anonymous):

got it thx c:

OpenStudy (haseeb96):

d.

OpenStudy (anonymous):

thanks! wanna help me with some others?

OpenStudy (anonymous):

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

OpenStudy (anonymous):

c